OpenStudy (blondisaurus):

Use log4^3= 0.7925 and log4^5 = 1.1610 Problem: log4^18

OpenStudy (blondisaurus):

I got to log4^3+log4^6 but 4 doesnt go into 6 so what do i do?

OpenStudy (raden):

set log4^18 = log {4^15*4^3} = log{(4^5)^3*4^3} = log(4^5)^3 + log(4^3) = 3*log(4^5) + log(4^3) = 3(....) + ..... = .....

OpenStudy (raden):

bla bla bla bla is given :)

OpenStudy (blondisaurus):

I think you did it wrong.

OpenStudy (blondisaurus):

What did you come out to..?

OpenStudy (raden):

which parts is it?

OpenStudy (raden):

tell me which part ?

OpenStudy (blondisaurus):

My friend solved it the way oyu explained and got 4. something but the answer is 2.085

OpenStudy (blondisaurus):

you* and i cant tell you what you did wrong because i dont understand it myself

OpenStudy (raden):

hahaha... it's not exactly. if u have told me wrong u have to know which parts i have wrong !!. if continue it like i have done before it should be : 3*log(4^5) + log(4^3) = 3(1.1610) + 0.7925 = 3.4830 + 0.7925 = 4.2755

OpenStudy (blondisaurus):

Im sorry Q.Q ill ask her

OpenStudy (raden):

ok, nevermind... :)